# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v1

@@ Abstract
To quantify how well theoretical predictions of structural ensembles agree with experimental measurements, we depend on the accuracy of forward models. These models are computational frameworks that generate observable quantities from molecular configurations based on empirical relationships linking specific molecular properties to experimental measurements. Bayesian Inference of Conformational Populations (BICePs) is a reweighting algorithm that reconciles simulated ensembles with ensemble-averaged experimental observations, even when such observations are sparse and/or noisy. This is achieved by sampling the posterior distribution of conformational populations under experimental restraints as well as sampling the posterior distribution of uncertainties due to random and systematic error. In this study, we enhance the algorithm for the refinement of empirical forward model (FM) parameters. We introduce and evaluate two novel methods for optimizing FM parameters. The first method treats FM parameters as nuisance parameters, integrating over them in the full posterior distribution. The second method employs variational minimization of a quantity called the BICePs score that reports the free energy of “turning on” the experimental restraints. This technique, coupled with improved likelihood functions for handling experimental outliers, facilitates force field validation and optimization, as illustrated in recent studies (Raddi et al. 2023, 2024). Using this approach, we refine parameters that modulate the Karplus relation, crucial for accurate predictions of $J$-coupling constants based on dihedral angles ($\phi$) between interacting nuclei. We validate this approach first with a toy model system, and then for human ubiquitin, predicting six sets of Karplus parameters for ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. This approach, which does not rely on any predetermined parameterization, enhances predictive accuracy and can be used for many applications.
@@ I Introduction
In the field of molecular modeling and dynamics, the accuracy of theoretical predictions that reflect real-world observations is crucial. Quantifying the agreement between theory and experiment is highly dependent on the accuracy of forward models—computational frameworks that predict observable quantities from molecular configurations. These models often depend on empirical relationships that link specific molecular properties to experimental measurements.
@@
Model validation and refinement of structural ensembles against NMR observables critically depends on reliable forward models (FMs) that have been robustly parameterized, so that FM error is minimal in the validation/refinement process. An important challenge in the parameterization of FMs is presented by random and systematic errors inherent to the experimental data. These errors need to be considered in the comparison and integration of experimental data with computational models for objective model selection and accurate uncertainty representation.
@@
A further challenge is presented by missing or insufficient examples of known structures than can be used to train forward models. For NMR observables that depend on backbone $\phi$-angles, such as $J$-coupling constants, the reference data from X-ray crystallography may be missing or dynamically averaged, creating large uncertainties in the correct $\phi$-angles. Numerous approaches Fröhlking, Bernetti, and Bussi (2023); Wang and Bax (1996); Habeck, Rieping, and Nilges (2005); Schmidt et al. (1999) have been developed to address some of these challenges. Some algorithms rely heavily on X-ray crystal structure data; others have many hyperparameters that need to be determined.
@@
To address these challenges, we extend the Bayesian Inference of Conformational Populations (BICePs) algorithmVoelz, Ge, and Raddi (2021); Raddi et al. (2023) to refine FM parameters. BICePs, a reweighting algorithm, refines structural ensembles against sparse and/or noisy experimental observables, and has been used in many previous applications.Voelz and Zhou (2014); Wan et al. (2020); Hurley et al. (2021); Raddi, Ge, and Voelz (2023) BICePs infers all possible sources of error by sampling the posterior distribution of these parameters directly from the data through MCMC sampling BICePs also computes a free energy-like quantity called the BICePs score that can be used for model selection and model parameterization.Ge and Voelz (2018); Raddi et al. (2023); Raddi and Voelz (2024)
@@
Recently, BICePs was enhanced with a replica-averaging forward model, making it a maximum-entropy (MaxEnt) reweighting method, and unique in that no adjustable regularization parameters are required to balance experimental information with the prior.Raddi et al. (2023) With this new approach, the BICePs score becomes a powerful objective function to parameterize optimal models. Here, we show that the BICePs score, which reflects the total evidence for a model, can be used for variational optimization of FM parameters. The BICePs score contains a form of inherent regularization, and has specialized likelihood functions that allow for the automatic detection and down-weighting of the importance of experimental observables subject to systematic error.Raddi et al. (2023)
To effectively refine FM parameters, we sample over the full posterior distribution of FM parameters. Through this approach, BICePs performs ensemble reweighting and FM parameter refinement simultaneously. Additionally, we show that by variational minimization of the BICePs score, we obtain the same result and show that the two approaches are equivalent, with each method requiring particular considerations. We first demonstrate our method’s effectiveness on a toy model system, and then optimize six distinct sets of Karplus parameters for the human protein ubiquitin, and compare our findings with previously established results. Through this, we aim to showcase a systematic and robust approach to enhancing the accuracy of theoretical predictions, thereby bridging the gap between computational models and experimental observations.
@@ II Theory
Posterior sampling of forward model parameters gives reliable parameter uncertainties.
@@
We now consider a specific forward model $g(X,\theta)$ with a set of FM parameters $\theta$ that we wish to additionally include in the posterior,
$p(X,\sigma,\theta|D)\propto p(D|X,\sigma,\theta)p(X)p(\sigma)p(\theta)$ (2)
@@
When BICePs is equipped with a replica-averaged forward model, it becomes a MaxEnt reweighting method in the limit of large numbers of replicas Pitera and Chodera (2012); Cavalli, Camilloni, and Vendruscolo (2013); Cesari, Reißer, and Bussi (2018); Roux and Weare (2013); Hummer and Köfinger (2015); Bonomi et al. (2016). Consider a set of $N$ replicas, $\mathbf{X}=\{X_{r}\}$, where $X_{r}$ is the conformational state being sampled by replica $r$. To compare the sampled replicas with ensemble-averaged experimental observables, we define a replica-averaged forward model $g(\mathbf{X},\theta)=\frac{1}{N}\sum^{N}_{r}g(X_{r},\theta)$. This quantity is an estimator of the true ensemble-average, with an error due to finite sampling for observable $j$ estimated using standard error of the mean (SEM):Pitera and Chodera (2012); Bonomi et al. (2016) $\sigma^{SEM}_{j}=\sqrt{\frac{1}{N}\sum\nolimits^{N}_{r}(g_{j}(X_{r},\theta)-\langle g_{j}(\mathbf{X},\theta)\rangle)^{2}}$. Thus, $\sigma^{SEM}_{j}$ decreases as the square root of the number of replicas.
In the scenario that observables can be collected into different types, e.g., a particular type of vicinal $J$-coupling, then each collection can be described with its own set of parameters and error distribution. For $K$ distinct sets of FM parameters $\mathbf{\theta}=\{\theta_{k}\}$, the joint posterior distribution for all parameters is
$p(X,\sigma,\theta|D)\propto\prod_{r=1}^{N}p(X_{r})\prod_{k=1}^{K}p(D_{k}|g(\mathbf{X},\theta_{k}),\sigma_{k})p(\sigma_{k})p(\theta_{k})$ (3)
where $\mathbf{X}$ is a set of $N$ conformation replicas, and $\theta_{k}$ is the $k^{th}$ set of FM parameters. The $k^{th}$ set has an uncertainty parameter $\sigma_{k}=\sqrt{(\sigma^{\text{SEM}}_{k})^{2}+(\sigma^{\text{B}}_{k})^{2}}$, that describes the total error, arising from both finite sampling $(\sigma^{\text{SEM}}_{k})^{2}$, and uncertainty in the experimental measurements, known as a Bayesian uncertainty parameter $\sigma^{\text{B}}_{k}$. The prior distribution of uncertainties $p(\sigma_{k})$ is treated as a non-informative Jeffrey’s prior $(\sigma_{k}^{-1})$ for each collection of observables, and the posterior of FM parameters $p(\theta|D)$ is recovered by marginalization over all $X$ and $\sigma$:
@@
In our methodology, Markov chain Monte Carlo (MCMC) is used to sample the posterior with acceptances following the Metropolis-Hastings (M-H) criterion. Our algorithm can be used with or without gradients. However, significantly faster convergence, especially in higher dimensions, is achieved through an integration of stochastic gradient descent approach. Our gradient descent approach allows for informed updates to the FM parameters, incorporating stochastic noise to facilitate the escape from local minima and enhance exploration of the parameter space.
@@
where $\theta_{\text{trial}}$ and $\theta_{\text{old}}$ denote the trial parameters and previous parameters, respectively. The learning rate is denoted by $l_{\text{rate}}$, $\nabla u$ signifies the computed gradient of BICePs energy function with respect to the parameters $\theta$, and $\eta$ scales the noise drawn from a standard normal distribution $\mathcal{N}(0,1)$.
This strategic parameter update protocol is designed to satisfy the M-H criterion, ensuring that each step in the parameter space not only moves towards minimizing the energy of the forward model but also adheres to the probabilistic acceptance of potentially non-optimal moves to avoid local optima traps. Ergodic sampling is ensured by "turning off" the gradient after burn-in. The sampling procedure involves: (1) acquiring derivatives of the FM parameters, (2) perturbing these parameters based on the derived information, (3) predict observables using perturbed FM parameters and compute the total energy, and (4) assessing the new energy against the previous to determine acceptance based on the M-H criterion. This ensures a thorough and effective search of the parameter space, leveraging both the landscape topology and stochastic elements to guide the exploration.
The Good-Bad model accounts for systematic error due to outlier measurements
BICePs now is equipped with sophisticated likelihood models that are extremely robust in the presence of systematic errorRaddi et al. (2023). Recently, we demonstrated the ability of the Student’s model to account for systematic error for force field optimizationRaddi and Voelz (2024). In this work, we use a likelihood function called the Good-Bad model to demonstrate the validity of forward model refinement. The derivatives of the Good-Bad model are far less complicated than the Student’s model.
The Good-Bad likelihood modelRaddi et al. (2023) assumes that the level of noise is mostly uniform, except for a few erratic measurements. This limits the number of uncertainty parameters that need to be sampled, while still capturing outliers. Consider a model where uncertainties $\sigma_{j}$ for particular observables $j$ are distributed about some typical uncertainty $\sigma^{B}$ according to a conditional probability $p(\sigma_{j}|\sigma^{B})$. We derive a posterior for the $k^{th}$ parameter set having a single uncertainty parameter $\sigma^{B}$ by marginalizing over all $\sigma_{j}$
@@
where $\sigma_{0}=\sqrt{(\sigma^{B})^{2}+(\sigma^{SEM})^{2}}$. Under the Good-Bad model, we say that the "good" data consists of observables normally distributed about their true values with effective variance $\sigma_{0}^{2}$, while the "bad" data is subject to systematic error, leading to a larger effective variance $\varphi^{2}\sigma_{0}^{2}$, where $\varphi\geq 1$. By this assignment, $p(\sigma_{j}|\sigma_{0})$ from equation 6 becomes
@@
$\begin{split}p(\mathbf{X},&\sigma_{0},\varphi,\theta_{k}|D)\propto\prod^{N}_{r=1}\left\{p(X_{r})\prod^{N_{d}}_{j=1}\int\limits_{0}^{1}d\omega\int\limits_{\sigma^{\text{SEM}}}^{\infty}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta_{k}))^{2}}{2\sigma_{j}^{2}}\right)\frac{\omega\delta(\sigma_{j}-\varphi\sigma_{0})+(1-\omega)\delta(\sigma_{j}-\sigma_{0})}{\sqrt{2\pi}\sigma_{j}}d\sigma_{j}\right\}\\ &=\prod^{N}_{r=1}\left\{p(X_{r})\prod^{N_{d}}_{j=1}\Biggl{(}\frac{\left(1-H\left(\sigma^{SEM}-\sigma_{0}\right)\right)}{2\sqrt{2\pi}\sigma_{0}}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta_{k}))^{2}}{2\sigma_{0}^{2}}\right)+\frac{\left(1-H\left(\sigma^{SEM}-\varphi\sigma_{0}\right)\right)}{2\varphi\sqrt{2\pi}\sigma_{0}}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta_{k}))^{2}}{2\varphi^{2}\sigma_{0}^{2}}\right)\Biggr{)}\right\},\end{split}$ (8)
@@
For a single set of FM parameters (for simplicity), the BICePs energy function, $u=-\log p(\mathbf{X},\sigma_{0},\varphi,\theta|D)$, the negative logarithm of the posterior in its full form is given by
@@
$\begin{split}u=\sum^{N}_{r=1}-\log(p(X_{r}))-N\sum^{N_{d}}_{j=1}\log\left[\frac{\left(1-H\left(\sigma^{SEM}-\sigma_{0}\right)\right)}{2\sqrt{2\pi}\sigma_{0}}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta))^{2}}{2\sigma_{0}^{2}}\right)+\frac{\left(1-H\left(\sigma^{SEM}-\varphi\sigma_{0}\right)\right)}{2\varphi\sqrt{2\pi}\sigma_{0}}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta))^{2}}{2\varphi^{2}\sigma_{0}^{2}}\right)\right],\end{split}$ (9)
and when $\varphi=1$ our energy function becomes
@@
$\begin{split}u=\sum^{N}_{r=1}-\log(p(X_{r}))+N\left[\sum_{j=1}^{N_{d}}-\log\left(\frac{1}{\sqrt{2\pi}\sigma_{j}}\right)+\frac{(d_{j}-g_{j}(\mathbf{X},\theta))^{2}}{2\sigma_{j}^{2}}-\log(p(\sigma_{j}))\right].\end{split}$ (10)
The first derivative of equation 9 with respect to the $i^{th}$ FM parameter $\theta_{i}$ is
$\frac{\partial u}{\partial\theta_{i}}=N\sum^{N_{d}}_{j=1}\frac{\partial g_{j}(X,\theta)}{\partial\theta_{i}}\frac{(d_{j}-g_{j}(X,\theta))}{\varphi^{2}\sigma_{0}^{2}}\frac{\Biggl{\{}\varphi^{3}\left(1-H\left(\sigma^{SEM}-\sigma_{0}\right)\right)\exp\left(\frac{\left(d_{j}-g_{j}(X,\theta)\right)^{2}}{2\varphi^{2}\sigma_{0}^{2}}\right)+\left(1-H\left(-\varphi\sigma_{0}+\sigma^{SEM}\right)\right)\exp\left(\frac{\left(d_{j}-g_{j}(X,\theta)\right)^{2}}{2\sigma_{0}^{2}}\right)\Biggr{\}}}{\Biggl{\{}\varphi\left(H\left(\sigma^{SEM}-\sigma_{0}\right)-1\right)\exp\left({\frac{\left(d_{j}-g_{j}(X,\theta)\right)^{2}}{2\varphi^{2}\sigma_{0}^{2}}}\right)+\left(H\left(-\varphi\sigma_{0}+\sigma^{SEM}\right)-1\right)\exp\left({\frac{\left(d_{j}-g_{j}(X,\theta)\right)^{2}}{2\sigma_{0}^{2}}}\right)\Biggr{\}}},\vspace{0.2cm}$ (11)
and in the case of $\varphi=1$ the gradient becomes
@@
$\frac{\partial u}{\partial\theta_{i}}=-N\left[\sum_{j=1}^{N_{d}}\frac{\partial g_{j}(X,\theta)}{\partial\theta_{i}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right].$ (12)

### v2

@@ Abstract
To quantify how well theoretical predictions of structural ensembles agree with experimental measurements, we depend on the accuracy of forward models. These models are computational frameworks that generate observable quantities from molecular configurations based on empirical relationships linking specific molecular properties to experimental measurements. Bayesian Inference of Conformational Populations (BICePs) is a reweighting algorithm that reconciles simulated ensembles with ensemble-averaged experimental observations, even when such observations are sparse and/or noisy. This is achieved by sampling the posterior distribution of conformational populations under experimental restraints as well as sampling the posterior distribution of uncertainties due to random and systematic error. In this study, we enhance the algorithm for the refinement of empirical forward model (FM) parameters. We introduce and evaluate two novel methods for optimizing FM parameters. The first method treats FM parameters as nuisance parameters, integrating over them in the full posterior distribution. The second method employs variational minimization of a quantity called the BICePs score that reports the free energy of “turning on” the experimental restraints. This technique, coupled with improved likelihood functions for handling experimental outliers, facilitates force field validation and optimization, as illustrated in recent studies (Raddi et al. 2023, 2024). Using this approach, we refine parameters that modulate the Karplus relation, crucial for accurate predictions of $J$-coupling constants based on dihedral angles ($\phi$) between interacting nuclei. We validate this approach first with a toy model system, and then for human ubiquitin, predicting six sets of Karplus parameters for ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. Finally, we demonstrate that our framework naturally generalizes optimization to any differentiable forward model, such as those constructed with neural networks. This approach, which does not rely on any predetermined parameterization, enhances predictive accuracy and can be used for many applications.
Keywords: Bayesian Inference | Conformational Populations | Forward Model Optimization | Karplus Relation | Neural Network-based Forward Models | Machine Learning for Structural Biology |
@@ I Introduction
In the field of molecular modeling and dynamics, the accuracy of theoretical predictions is crucial. The quantitative agreement between theory and experiment is typically characterized through the use of accurate forward models—computational frameworks that predict observable quantities from molecular configurations. These models often depend on empirical relationships that link specific molecular properties to experimental measurements.
@@
Model validation and refinement of structural ensembles against NMR observables critically depends on reliable forward models (FMs) that have been robustly parameterized to minimize FM error. In refining these FMs, it is important to consider random and systematic errors inherent to the experimental data, which in turn contribute to overall uncertainty of integrative models. It is usually challenging to obtain complete knowledge of such errors; instead, such errors typically need to be inferred from the data.
@@
A further challenge is presented by missing or insufficient examples of known structures that can be used to train forward models. For NMR observables that depend on backbone $\phi$-angles, such as $J$-coupling constants, the reference data from X-ray crystallography may be missing or dynamically averaged, creating large uncertainties in the correct $\phi$-angles. Numerous approaches have been developed to address some of these challenges. Fröhlking, Bernetti, and Bussi (2023); Wang and Bax (1996); Habeck, Rieping, and Nilges (2005); Schmidt et al. (1999) Some algorithms rely heavily on X-ray crystal structure data; others have many hyperparameters that need to be determined.
@@
To address these challenges, we extend the Bayesian Inference of Conformational Populations (BICePs) algorithmVoelz, Ge, and Raddi (2021); Raddi et al. (2025) to refine FM parameters. BICePs isa reweighting algorithm to refine structural ensembles against sparse and/or noisy experimental observables, which has been used in many previous applications.Voelz and Zhou (2014); Wan et al. (2020); Hurley et al. (2021); Raddi, Ge, and Voelz (2023) BICePs infers all possible sources of error by sampling the posterior distribution of these parameters directly from the data through MCMC sampling. BICePs also computes a free energy-like quantity called the BICePs score that can be used for model selection and model parameterization.Ge and Voelz (2018); Raddi et al. (2025); Raddi and Voelz (2024)
@@
Recently, BICePs was enhanced with a replica-averaging forward model, making it a maximum-entropy (MaxEnt) reweighting method, and unique in that no adjustable regularization parameters are required to balance experimental information with the prior.Raddi et al. (2025) With this new approach, the BICePs score becomes a powerful objective function to parameterize optimal models. Here, we show that the BICePs score, which reflects the total evidence for a model, can be used for variational optimization of FM parameters. The BICePs score contains a form of inherent regularization, and is particularly powerful when used with specialized likelihood functions that allow for the automatic detection and down-weighting of the importance of experimental observables subject to systematic error.Raddi et al. (2025)
To effectively refine FM parameters, we use BICePs to sample over the entire posterior distribution of FM parameters, enabling simultaneous ensemble reweighting and parameter refinement. Additionally, we show that by variational minimization of the BICePs score, we obtain the same result and show that the two approaches are equivalent, with each method requiring particular considerations. We first demonstrate our method’s effectiveness using a toy model and then apply it to optimize six distinct sets of Karplus parameters for the human protein ubiquitin. As a proof-of-concept for more complex models, we also train a neural network to learn $J$-coupling constants from dihedral angles using the BICePs score as a loss function. We then discuss how this idea could be applied to neural networks for chemical shift prediction.Han et al. (2024); Yang, Chakraborty, and White (2021); Meiler (2003); Shen and Bax (2010) Of course, our framework naturally extends to other empirical forward models with tunable parameters—for example, hydrogen-deuterium exchange (HDX) models with adjustable hydrogen-bonding and burial weights, or paramagentic relaxation enhancement (PRE) models with system-specific correlation times and scaling factors. These examples illustrate the generality and flexibility of our approach for refining a wide range of forward models to improve agreement between simulations and experiment.
@@ II Theory
In this section, we develop two different theoretical approaches to optimize forward models. In the first method, we include forward model parameters in the posterior, and sample the full posterior distribution of all parameters. In the second approach, we optimize the forward model by minimizing the BICePs score, a free energy-like quantity that can be used for variational optimization.
II.1 Forward model optimization by posterior sampling of forward model parameters
@@
For a specific forward model $g(X,\theta)$ with a set of FM parameters $\theta=(\theta_{1},\theta_{2},...,\theta_{m})$, where $m$ is the number of parameters, we can additionally include the parameters in the posterior,
$p(X,\sigma,\theta|D)\propto p(D|X,\sigma,\theta)p(X)p(\sigma)p(\theta).$ (2)
@@
When BICePs is used witha replica-averaged forward model, it becomes a MaxEnt reweighting method in the limit of large numbers of replicas Pitera and Chodera (2012); Cavalli, Camilloni, and Vendruscolo (2013); Cesari, Reißer, and Bussi (2018); Roux and Weare (2013); Hummer and Köfinger (2015); Bonomi et al. (2016). Consider a set of $N$ replicas, $\mathbf{X}=\{X_{r}\}$, where $X_{r}$ is the conformational state being sampled by replica $r$. To compare the sampled replicas with ensemble-averaged experimental observables, we define a replica-averaged forward model $g(\mathbf{X},\theta)=\frac{1}{N}\sum^{N}_{r}g(X_{r},\theta)$. This quantity is an estimator of the true ensemble average, with an error due to finite sampling for the observable $j$ estimated using the standard error of the mean (SEM):Pitera and Chodera (2012); Bonomi et al. (2016) $\sigma^{\text{SEM}}_{j}=\sqrt{\frac{1}{N}\sum\nolimits^{N}_{r}(g_{j}(X_{r},\theta)-\langle g_{j}(\mathbf{X},\theta)\rangle)^{2}}$. Thus, $\sigma^{\text{SEM}}_{j}$ decreases as the square root of the number of replicas.
For a single forward model with parameters $\theta$, the joint posterior distribution for all parameters is
$p(X,\sigma,\theta|D)\propto\prod_{r=1}^{N}p(X_{r})p(D|g(\mathbf{X},\theta),\sigma)p(\sigma)p(\theta)$ (3)
where $\mathbf{X}$ is a set of $N$ conformation replicas. In many cases, we wish to optimize several forward models at once. For example, there are unique sets of Karplus parameters in the forward models for vicinal $J$-coupling constant observables ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, etc. If there are $K$ different forward models $\theta^{(1)},\theta^{(2)},...\theta^{(K)}$, each with their own set of parameters and error distributions, we can describe their full joint posterior distribution as
$\begin{split}&p(X,\sigma,\theta^{(1)},\theta^{(2)},...\theta^{(K)}|D)\\ &\propto\prod_{r=1}^{N}p(X_{r})\prod_{k=1}^{K}p(D_{k}|g(\mathbf{X},\theta^{(k)}),\sigma_{k})p(\sigma_{k})p(\theta^{(k)})\end{split}$ (4)
where $D_{k}$ are the experimental observables predicted by the $k^{th}$ forward model, $\theta^{(k)}$ are the parameters for the $k^{th}$ FM, and $\sigma_{k}=\sqrt{(\sigma^{\text{SEM}}_{k})^{2}+(\sigma^{\text{B}}_{k})^{2}}$ describing the total error for the $k^{th}$ observable. The total error originates from both finite sampling through the standard error of the mean $(\sigma^{\text{SEM}}_{k})$, and uncertainty in the experimental measurements, known as a Bayesian uncertainty parameter $\sigma^{\text{B}}_{k}$. The prior distribution of uncertainties $p(\sigma_{k})$ is treated as a non-informative Jeffrey’s prior $(\sigma_{k}^{-1})$ for each collection of observables, and the posterior of FM parameters $p(\theta|D)$ is recovered by marginalization over all $X$ and $\sigma$:
@@
In our methodology, Markov chain Monte Carlo (MCMC) is used to sample the posterior with acceptances following the Metropolis-Hastings criterion. Our algorithm can be used with or without gradients. However, significantly faster convergence, especially in higher dimensions, is achieved through an integration of stochastic gradient descent approach. Our gradient descent approach allows for informed updates to the FM parameters, incorporating stochastic noise to facilitate the escape from local minima and enhance exploration of the parameter space.
@@
where $\theta_{\text{trial}}$ and $\theta_{\text{old}}$ denote the trial parameters and previous parameters, respectively. The learning rate is denoted by $l_{\text{rate}}$, $\nabla u$ signifies the calculated gradient of BICePs energy function $u=-\ln p(X,\sigma,\theta|D)$ with respect to the parameters $\theta$, and $\eta$ scales the noise drawn from a standard normal distribution $\mathcal{N}(0,1)$.
This parameter update strategy is designed to satisfy the Metropolis-Hastings criterion, ensuring that each step in the parameter space not only moves towards minimizing the energy of the forward model but also adheres to the probabilistic acceptance of potentially non-optimal moves to avoid local optima traps. Ergodic sampling is ensured by "turning off" the gradient after burn-in. The sampling procedure involves: (1) acquiring derivatives of the FM parameters, (2) perturbing these parameters based on the derived information, (3) predicting observables using perturbed FM parameters and computing the total energy, and (4) assessing the new energy against the previous to determine acceptance based on the Metropolis-Hastings criterion. This ensures a thorough and effective search of the parameter space, taking advantage of both the landscape topology and stochastic elements to guide the exploration.
The Good-Bad model accounts for systematic error due to outlier measurements.
Recent work has developed sophisticated likelihood models for BICePs, designed to detect outliers, that are robust in the presence of systematic errorRaddi et al. (2025). Raddi et al. (2025) developed two likelihood models for this purpose: the Student’s model and the Good-Bad model. In related work using BICePs to perform automated force field optimization, we used the Student’s model.Raddi and Voelz (2024) Here, in this work, we use the Good-Bad model, primarily because the derivatives of the Good-Bad model are less complicated than the Student’s model.
The Good-Bad likelihood modelRaddi et al. (2025) assumes that the level of noise is mostly uniform, except for a few erratic measurements. This limits the number of uncertainty parameters that need to be sampled, while still capturing outliers. Consider a model in which the uncertainties $\sigma_{j}$ for particular observables $j$ are distributed with respect to some typical uncertainty $\sigma^{B}$ according to a conditional probability $p(\sigma_{j}|\sigma^{B})$. We derive a posterior for the parameters of the $k^{th}$ FM having a single uncertainty parameter $\sigma^{B}$ by marginalizing over all $\sigma_{j}$,
@@
where $\sigma_{0}=\sqrt{(\sigma^{B})^{2}+(\sigma^{SEM})^{2}}$. Under the Good-Bad model, we say that the "good" data consist of observables normally distributed about their true values with effective variance $\sigma_{0}^{2}$, while the "bad" data are subject to systematic error, leading to a larger effective variance $\varphi^{2}\sigma_{0}^{2}$, where $\varphi\geq 1$.
By this assignment, $p(\sigma_{j}|\sigma_{0})$ from equation 7 becomes
@@
[…]
@@
For a single set of FM parameters (for simplicity), the BICePs energy function, $u=-\log p(\mathbf{X},\sigma_{0},\varphi,\theta^{(k)}|D)$, the negative logarithm of the posterior in its full form is given by
@@
$\begin{split}u=\sum^{N}_{r=1}-\log(p(X_{r}))-N\sum^{N_{d}}_{j=1}\log\left[\frac{1-H(\sigma^{SEM}-\sigma_{0})}{2}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\sigma_{0}^{2}\right)+\frac{1-H(\sigma^{SEM}-\varphi\sigma_{0})}{2\varphi}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\varphi^{2}\sigma_{0}^{2}\right)\right].\end{split}$ (10)
When $\varphi=1$, this energy function is the same the result we would get if we used a Gaussian likelihood function,
@@
$\begin{split}u=\sum^{N}_{r=1}-\log(p(X_{r}))+N\left[\sum_{j=1}^{N_{d}}-\log\left(\frac{1}{\sqrt{2\pi}\sigma_{0}}\right)+\frac{(d_{j}-g_{j}(\mathbf{X},\theta))^{2}}{2\sigma_{0}^{2}}-\log(p(\sigma_{0}))\right].\end{split}$ (11)
The first derivative of equation 10 with respect to the $i^{th}$ FM parameter in the parameter set $\theta^{(k)}$ is
@@
$\frac{\partial u}{\partial\theta^{(k)}_{i}}=N\sum_{j=1}^{N_{d}}\frac{\partial g_{j}(\mathbf{X},\theta^{(k)})}{\partial\theta^{(k)}_{i}}\,\frac{(d_{j}-g_{j}(\mathbf{X},\theta^{(k)}))}{\sigma_{0}^{2}}\,\left(\frac{\varphi^{2}\left(1-H\left(\sigma^{SEM}-\sigma_{0}\right)\right)\mathcal{N}_{\varphi^{2}}^{-1}+\left(1-H\left(-\varphi\sigma_{0}+\sigma^{SEM}\right)\right)\mathcal{N}_{1}^{-1}}{\varphi^{2}\left(H\left(\sigma^{SEM}-\sigma_{0}\right)-1\right)\mathcal{N}_{\varphi^{2}}^{-1}+\varphi^{2}\left(H\left(-\varphi\sigma_{0}+\sigma^{SEM}\right)-1\right)\mathcal{N}_{1}^{-1}}\right),$ (12)
where $\mathcal{N}_{1}^{-1}=\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\sigma_{0}^{2}\right)^{-1}$ and $\mathcal{N}_{\varphi^{2}}^{-1}=\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\varphi^{2}\sigma_{0}^{2}\right)^{-1}$. Both represent the reciprocal of Gaussian likelihoods, where $\mathcal{N}_{1}^{-1}$ and $\mathcal{N}_{\varphi^{2}}^{-1}$ denote inverse-weighted forms of normal distributions with variances $\sigma_{0}^{2}$ and $\varphi^{2}\sigma_{0}^{2}$, respectively. When $\varphi=1$, the gradient becomes
$\frac{\partial u}{\partial\theta^{(k)}_{i}}=-N\sum_{j=1}^{N_{d}}\frac{\partial g_{j}(\mathbf{X},\theta^{(k)})}{\partial\theta^{(k)}_{i}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta^{(k)}))}{\sigma_{0}^{2}}.$ (13)